# Understanding the Ranking Loss for Recommendation with Sparse User Feedback

### v1

@@ Abstract.
Click-through rate (CTR) prediction holds significant importance in the realm of online advertising. While many existing approaches treat it as a binary classification problem and utilize binary cross entropy (BCE) as the optimization objective, recent advancements have indicated that combining BCE loss with ranking loss yields substantial performance improvements. However, the full efficacy of this combination loss remains incompletely understood. In this paper, we uncover a new challenge associated with BCE loss in scenarios with sparse positive feedback, such as CTR prediction: the gradient vanishing for negative samples. Subsequently, we introduce a novel perspective on the effectiveness of ranking loss in CTR prediction, highlighting its ability to generate larger gradients on negative samples, thereby mitigating their optimization issues and resulting in improved classification ability. Our perspective is supported by extensive theoretical analysis and empirical evaluation conducted on publicly available datasets. Furthermore, we successfully deployed the ranking loss in Tencent’s online advertising system, achieving notable lifts of 0.70% and 1.26% in Gross Merchandise Value (GMV) for two main scenarios. The code for our approach is openly accessible at the following GitHub repository: https://github.com/SkylerLinn/Understanding-the-Ranking-Loss.
copyright: acmcopyrightjournalyear: 2024doi: XXXXXXX.XXXXXXX11footnotetext: Equal contribution22footnotetext: Corresponding author
@@ 1. Introduction
(a) Training/Validation LogLoss
(b) Loss Landscapes
Figure 1. (a) Training/validation LogLoss dynamic of BCE method (blue) and Combined-Pair (red). (b) Loss landscape of BCE method and Combined-Pair.
@@
In recent decades, users have encountered abundant information while navigating websites or mobile applications. This inundation presents significant challenges for electronic retailers, content providers, and online advertising platforms as they strive to recommend appropriate items to individual users within specific contexts. Thus, the deployment of recommendation systems has become widespread, enabling the prediction of users’ preferences from a vast pool of candidate items. For instance, in the domain of cost-per-click (CPC) advertising, advertising platforms must bid for each advertisement based on the estimated value of the impression, which relies on the bid value and the estimated Click-through rate (CTR). Consequently, accurately predicting user response has emerged as a critical factor, attracting substantial research attention.
@@
CTR prediction (Chen et al., 2020; McMahan et al., 2013; Rendle, 2010a; He et al., 2014) is usually formulated as a binary classification problem and optimized by a binary cross entropy (BCE) loss. Some recent works (Li et al., 2015; Yan et al., 2022; Sheng et al., 2023; Bai et al., 2023) in the industry propose to combine the BCE loss with the ranking loss, which was usually used in Learning to Rank (LTR) (Burges et al., 2005, 2006; Cao et al., 2007; Zheng et al., 2007; Liu et al., 2009). For example, Combined-Pair (Li et al., 2015) stands as one of the pioneering attempts to combine a pairwise loss with a pointwise loss for CTR prediction in the Twitter Timeline. Yan et al. and  Bai et al. proposed to combine regression loss with ranking loss to better trade-off between the regression and ranking objectives.  Sheng et al. proposed to employ two logits to jointly optimize ranking and calibration objectives. Such combination loss is widely adopted in real-world recommendation systems in Twitter (Li et al., 2015), Google (Yan et al., 2022; Bai et al., 2023), and Alibaba (Sheng et al., 2023).
Existing literature primarily attributes the success of combining classification and ranking losses to the enhancement of ranking ability (Li et al., 2015; Sheng et al., 2023). These studies validate their hypothesis by observing an improvement in the Area Under the Curve (AUC), a metric used to assess ranking quality, on the testing data. However, we are intrigued by how the combination loss impacts the model’s primary optimization objective: the classification ability, as measured by the LogLoss. Surprisingly, our observations reveal a reduction in LogLoss on the testing data with sparse positive feedback, as depicted in Fig. 1(a). This implies that incorporating an additional ranking loss not only enhances the model’s ranking ability but also improves its classification ability.
@@
To further investigate the cause of this improvement, we delve into two potential factors: the model’s generalization capability and easier optimization procedure during training. We monitor the LogLoss during training using the combination method. To our surprise, we find that it is significantly reduced compared to the method that solely employs Binary Cross-Entropy (BCE) loss, referred to as the BCE method. Moreover, we visualize the loss landscape of Combined-Pair and the BCE method, and observe a notable distinction. The BCE method exhibits a relatively flat landscape, whereas the Combined-Pair method demonstrates a significantly steeper landscape, as illustrated in Figure 1(b). This observation suggests that the BCE method is susceptible to gradient vanishing issues.
@@
We proceed to investigate the gradients of the Binary Cross Entropy (BCE) loss and the BCE-ranking combination loss. In scenarios with sparse positive feedback, such as ad Click-Through Rate (CTR) prediction where only a small fraction of samples are positive (clicks), negative samples tend to receive small gradients when using the BCE loss, leading to optimization challenges. However, in the combination loss, the ranking loss contributes significantly larger gradients, effectively mitigating the aforementioned gradient vanishing problem.
@@
To further validate our perspective, we conduct comprehensive experiments. Firstly, we generate artificial datasets with varying degrees of positive sample sparsity and observe that the sparser the positive samples, the greater the performance improvement achieved by incorporating a ranking loss alongside the BCE loss. Secondly, in addition to the combination loss, we explore alternative approaches that address the gradient vanishing issue, such as Focal Loss (Lin et al., 2017).
Lastly, we successfully deployed the classification-ranking combination loss in the CTR prediction within Tencent’s online advertising system, resulting in substantial revenue increases. In summary, our contributions can be summarized as follows:
@@
  • We uncover a novel challenge associated with binary cross entropy loss in recommendation scenarios with sparse positive feedback: the gradient vanishing of negative samples.
@@
  • We present a fresh perspective to comprehend the effectiveness of the classification-ranking combination loss in recommendation systems. The inclusion of ranking loss introduces larger gradients for negative samples, effectively addressing their optimization challenges during training.
  • We substantiate our claims through theoretical analysis, offline experiments, and online empirical evaluations. Our findings demonstrate that the ranking loss generates significantly larger and compatible gradients for negative samples, resulting in improved optimization during training and enhanced classification performance on the testing data.
2. Related Work and Preliminaries
@@ 2.1. Click-Through Rate Prediction
Current research in the CTR prediction task primarily focuses on model structure and optimization objectives. Regarding model structure, recent studies have primarily aimed to explore ways to capture higher-order interaction information effectively. This has led to the proposal of various model structures such as Factorization Machines (FM) family (Rendle, 2010b; Juan et al., 2017; Pan et al., 2018; Sun et al., 2021), Wide & Deep (Wang et al., 2017), DeepFM (Guo et al., 2017), IPNN (Qu et al., 2016), xDeepFM (Lian et al., 2018), DCN V2 (Wang et al., 2021), Final MLP (Mao et al., 2023) and Multi-Embedding (Guo et al., 2023a). Three main paradigms have been introduced for optimization objectives: pointwise, pairwise, and listwise approaches, which will be discussed in the following sections.
@@ 2.2. Binary Cross Entropy
(1) $\mathcal{L}_{\text{BCE}}=-\frac{1}{N}\sum_{i\in N}[y_{i}\log(\sigma(z_{i}))+(1-y_{i})\log(1-\sigma(z_{i}))],$
where $N$ denotes the number of samples, $z_{i}$ represents the logit of $i$-th sample, and $y_{i}$ denotes the corresponding binary label, 1 for click, and 0 for non-click.
@@ 2.3. Learning to Rank
(2) $\mathcal{L}_{RankNet}=-\frac{1}{N^{2}}\sum\limits_{i,j\in N}$ $[y_{ij}\log(\sigma(z_{i}-z_{j}))$
@@
where $y_{ij}\in\{0,0.5,1\}$, corresponding to the conditions that $y_{i}<y_{j},y_{i}=y_{j},y_{i}>y_{j}$, respectively. Following enhancements in the optimization process (Burges et al., 2006) and the incorporation of hinge loss (Tagami et al., 2013) have led to further improvements in performance.
Listwise methods encourage positive samples to have high rankings within the list of all samples. For example, ListNet (Cao et al., 2007) defines its loss as:
(3) $\mathcal{L}_{ListNet}=-\frac{1}{N}\sum\limits_{i\in N}\log\frac{exp[z_{i}]}{\sum\limits_{k\in N}exp[z_{k}]}.$
Some other listwise approaches have also achieved remarkable results, such as ListMLE (Xia et al., 2008), BayesRank (Kuo et al., 2009) and PiRank (Swezey et al., 2021). Recently, studies have revealed that such methods lack calibration capabilities, potentially leading to training instability Consequently, improved approaches like CalSoftmax (Yan et al., 2022) have been proposed.
@@ 2.4. Tailor Ranking Loss into CTR Prediction
In the context of online advertising, recent studies (Sheng et al., 2023; Yan et al., 2022; Bai et al., 2023) argued that it is not easy to achieve decent overall outcomes solely by relying on a single form of the objective function. Consequently, several research that combine classification loss and ranking loss have been proposed with the following objective paradigm:
@@
(5) $\mathcal{L}_{rank}^{RCR}=-\frac{1}{N}\sum\limits_{i\in N}\log\frac{\sigma(z_{i})}{\sum\limits_{k\in N}\sigma(z_{k})},$
@@
(6) $\mathcal{L}_{clf}^{JRC}=-\frac{1}{N}\sum\limits_{i\in N}\log\frac{exp[z_{iy_{i}}]}{exp[z_{i0}]+exp[z_{i1}]},$
(7) $\mathcal{L}_{rank}^{JRC}=-\frac{1}{N}\sum\limits_{i\in N}\log\frac{exp[z_{iy_{i}}]}{\sum\limits_{k\in N}exp[z_{ky_{i}}]},$
where $z_{i1}$ and $z_{i0}$ are the click/non-click logits of the $i$-th sample, respectively. $z_{ky_{i}}$ is the click logit of $k$-th sample if $i$-th samples is positive. Otherwise, $z_{ky_{i}}$ signifies the non-click logit. In this section 3 and section 4, we will conduct detailed empirical experiments and analyses on these methods.
3. Understanding Ranking Loss from Gradient Vanishing Perspective
The classification-ranking combination methods are widely used in real-world recommendation systems, e.g., Twitter (Li et al., 2015), Google (Yan et al., 2022; Bai et al., 2023), and Alibaba (Sheng et al., 2023). However, the effectiveness of such methods remains elusive, which drives us to explore the underlying mechanisms at play.
In our following study, we utilize Combined-Pair loss (Li et al., 2015) as a representative classification-ranking combination loss, using DCNv2 (Wang et al., 2021) as the model architecture. Combined-Pair is defined as a combination of a binary cross entropy (BCE) loss and a RankNet loss (Burges et al., 2005), without the terms of $y_{i}=y_{j}$, defined as:
(8) $\mathcal{L}^{CP}$ $=\alpha\mathcal{L}_{\text{BCE}}+(1-\alpha)\mathcal{L}_{\text{rank}}^{CP}$
$=$ $-[\alpha\cdot\frac{1}{N}\sum_{i\in N}[y_{i}\log(\sigma(z_{i}))+(1-y_{i})\log(1-\sigma(z_{i}))]$
$+(1-\alpha)\cdot\frac{1}{N_{+}N_{-}}\sum\limits_{i\in N_{+}}\sum\limits_{j\in N_{-}}\log(\sigma(z_{i}^{(+)}-z_{j}^{(-)}))],$
where $N_{+}$ and $N_{-}$ represent the numbers of positive and negative samples, respectively. In the following discussion, we name the model that optimizes the Combined-Pair loss as Combined-Pair method, or Combined-Pair in short, and the model that optimizes only the binary cross entropy (BCE) loss as BCE method.
@@ 3.1. Investigation of Classification Ability
Existing work mainly attributes the success of classification-ranking combination loss to enhancing the ranking ability (Li et al., 2015; Sheng et al., 2023). However, we are curious about its impact on the main objective of CTR prediction: the classification capability. We monitor the LogLoss of the BCE method and Combined-Pair on the testing data of the public Criteo dataset, as the dashed lines in Fig. 1(a). To our surprise, we have the following findings:
Finding 1. Combined-Pair gets a significantly lower LogLoss than BCE method on the testing data, indicating that it improves the classification capability by involving the additional ranking loss, rather than only the ranking ability.
The improvement in the classification loss can be caused by either better generalization or easier model training when adding the ranking loss. To figure it out, we also monitor the LogLoss of both methods on the training set of Criteo (Labs, 2014), as the solid lines shown in Fig. 1(a). We have a following interesting finding:
Finding 2. Combined Pair gets a significantly lower LogLoss than the BCE method on the training data, indicating that it helps optimize the BCE loss.
Besides, we analyze the disparities between the loss landscapes of Combined-Pair and the BCE method. As shown in Fig. 1(b), we observe that the loss landscape of the BCE method exhibits flatter than Combined-Pair, suggesting a potential issue of gradient vanishing. We then delve further into the reason behind these observations by investigating the gradients of these two losses.

### v2

@@ Abstract.
Click-through rate (CTR) prediction is a crucial area of research in online advertising. While binary cross entropy (BCE) has been widely used as the optimization objective for treating CTR prediction as a binary classification problem, recent advancements have shown that combining BCE loss with an auxiliary ranking loss can significantly improve performance. However, the full effectiveness of this combination loss is not yet fully understood. In this paper, we uncover a new challenge associated with the BCE loss in scenarios where positive feedback is sparse: the issue of gradient vanishing for negative samples. We introduce a novel perspective on the effectiveness of the auxiliary ranking loss in CTR prediction: it generates larger gradients on negative samples, thereby mitigating the optimization difficulties when using the BCE loss only and resulting in improved classification ability. To validate our perspective, we conduct theoretical analysis and extensive empirical evaluations on public datasets. Additionally, we successfully integrate the ranking loss into Tencent’s online advertising system, achieving notable lifts of 0.70% and 1.26% in Gross Merchandise Value (GMV) for two main scenarios. The code is openly accessible at: https://github.com/SkylerLinn/Understanding-the-Ranking-Loss.
Recommendation Systems; CTR Prediction; Gradient Vanishing; Ranking Loss
copyright: acmcopyrightjournalyear: 2024copyright: rightsretainedconference: Proceedings of the 30th ACM SIGKDD Conference on Knowledge Discovery and Data Mining; August 25–29, 2024; Barcelona, Spainbooktitle: Proceedings of the 30th ACM SIGKDD Conference on Knowledge Discovery and Data Mining (KDD ’24), August 25–29, 2024, Barcelona, Spaindoi: 10.1145/3637528.3671565isbn: 979-8-4007-0490-1/24/08ccs: Information systems Display advertisingccs: Computing methodologies Neural networksccs: Computing methodologies Factorization methods
@@ 1. Introduction
In recent decades, users have encountered abundant information while navigating websites or mobile applications. This inundation presents significant challenges for electronic retailers, content providers, and online advertising platforms as they strive to recommend appropriate items to individual users within specific contexts. Thus, the deployment of recommendation systems has become widespread, enabling the prediction of users’ preferences from a vast pool of candidate items.
For instance, in effective cost per mille (eCPM) advertising, advertising platforms must bid for each advertisement based on the estimated value of the impression, which relies on the bid value and the estimated Click-through rate (CTR). Consequently, accurately predicting user response has emerged as a critical factor, attracting substantial research attention.
@@
CTR prediction (Chen et al., 2020; McMahan et al., 2013; Rendle, 2010a; He et al., 2014) is usually formulated as a binary classification problem and optimized by a binary cross entropy (BCE) loss. Some recent works (Li et al., 2015; Yan et al., 2022; Sheng et al., 2023; Bai et al., 2023) in the industry propose to combine the BCE loss with an auxiliary ranking loss, which was usually used in Learning to Rank (LTR) (Burges et al., 2005, 2006; Cao et al., 2007; Zheng et al., 2007; Liu et al., 2009). For example, Combined-Pair (Li et al., 2015) stands as one of the pioneering attempts to combine a pairwise loss with a pointwise loss for CTR prediction in the Twitter Timeline. Yan et al. and  Bai et al. proposed to combine regression loss with ranking loss to better trade-off between the regression and ranking objectives (Yan et al., 2022; Bai et al., 2023).  Sheng et al. proposed to employ two logits to optimize ranking and calibration objectives (Sheng et al., 2023) jointly. Such combination loss is widely adopted in real-world recommendation systems in Twitter (Li et al., 2015), Google (Yan et al., 2022; Bai et al., 2023), and Alibaba (Sheng et al., 2023).
The prevailing literature predominantly attributes the success of combining classification and ranking losses to the augmentation of ranking capability (Li et al., 2015; Sheng et al., 2023). These studies substantiate their hypothesis by observing an enhancement in the Area Under the Curve (AUC), a metric commonly employed to evaluate ranking quality. However, our curiosity lies in investigating the impact of the combination loss on the model’s primary optimization objective: the classification ability, as measured by the BCE loss metric. To explore this, we conducted a comparative analysis between the BCE method, which solely employs the Binary Cross-Entropy (BCE) loss, and the Combined-Pair approach, which incorporates a BCE-ranking combination loss. Our evaluation was performed on the Criteo dataset, with artificial weights on positive samples to simulate the sparse positive scenario.
Surprisingly, our findings on the validation set, as depicted in Fig. 1(a), revealed a reduction in the BCE loss of the Combined-Pair method (the red dashed line) compared to that of the BCE method (the blue dashed line). This intriguing observation suggests that the inclusion of an additional ranking loss not only enhances the model’s ranking ability but also improves its classification ability.
@@
To investigate the cause of this improvement further, we delve into the optimization procedure during model training. In particular, we monitor the BCE loss of these two methods during model training, as the solid lines in Fig. 1(a). To our surprise, we find that the BCE loss of the Combined-Pair (the red solid line) experiences a significant reduction compared to that of the BCE method (the blue solid line). Moreover, we visualize the loss landscape of these two methods using Contour Plots & Random Directions (Goodfellow et al., 2014). Our analysis reveals that the BCE method exhibits a relatively flat landscape, indicating a slower optimization process. In contrast, the Combined-Pair method demonstrates a significantly steeper landscape, as illustrated in Fig. 1(b).
(a) BCE Loss Dynamics along epochs.
(b) Loss Landscapes.
Figure 1. (a) BCE Loss Dynamics along epochs on the training and validation set and (b) Loss Landscape of BCE method (blue) and Combined-Pair (red).
@@
We proceed to investigate the gradients of the BCE method and the Combined-Pair. In scenarios with sparse positive feedback, such as ad CTR prediction where only a small fraction of samples are positive (clicks), we demonstrate that negative samples get small gradients in the BCE method, leading to optimization difficulties. However, the ranking loss in the Combined-Pair contributes significantly larger gradients, effectively mitigating the gradient vanishing problem.
@@
To further validate our perspective, we conduct comprehensive experiments. Firstly, we generate artificial datasets with varying degrees of positive sample sparsity and observe that the sparser the positive samples, the greater the performance improvement achieved by incorporating an auxiliary ranking loss. Secondly, in addition to the classification-ranking combination loss, we explore alternative approaches that address the gradient vanishing issue, such as Focal Loss (Lin et al., 2017) and a new Combined-Contrastive method. Lastly, we successfully deployed the classification-ranking combination loss in the CTR prediction within Tencent’s online advertising system, resulting in substantial revenue increases. In summary, our contributions can be summarized as follows:
@@
  • We uncover a challenge associated with binary cross entropy loss in recommendation scenarios with sparse positive feedback: the gradient vanishing of negative samples.
@@
  • We present a novel perspective on the effectiveness of involving an auxiliary ranking loss in recommendation systems: it introduces larger gradients for negative samples, addressing the gradient vanishing issue.
  • We substantiate our claims through theoretical analysis, offline experiments, and online empirical evaluations.
2. Related Work
@@ 2.1. Click-Through Rate Prediction
Current research in the CTR prediction task primarily focuses on model structure and optimization objectives. Regarding model structure, recent studies have primarily aimed to explore ways to capture higher-order interaction information effectively. This has led to the proposal of various model structures such as Factorization Machines (FM) family (Rendle, 2010b; Juan et al., 2017; Pan et al., 2018a; Sun et al., 2021), Wide&Deep (Wang et al., 2017), DeepFM (Guo et al., 2017), IPNN (Qu et al., 2016), xDeepFM (Lian et al., 2018), DCN V2 (Wang et al., 2021), Final MLP (Mao et al., 2023) and Multi-Embedding (Guo et al., 2024). As for optimization objectives, three main loss paradigms have been introduced: pointwise, pairwise, and listwise approaches, which will be discussed in the following sections.
@@ 2.2. Binary Cross Entropy
(1) $\mathcal{L}_{\text{BCE}}=-\frac{1}{N}\sum_{i=1}^{N}[y_{i}\log(\sigma(z_{i}))+(1-y_{i})\log(1-\sigma(z_{i}))],$
where $N$ denotes the number of samples, $z_{i}$ represents the logit of $i$-th sample, and $y_{i}$ denotes the corresponding binary label, 1 for click and 0 for non-click.
@@ 2.3. Learning to Rank
(2) $\mathcal{L}_{\text{RankNet}}=-\frac{1}{N^{2}}\sum\limits_{i=1}^{N}\sum\limits_{j=1}^{N}$ $[y_{ij}\log(\sigma(z_{i}-z_{j}))$
@@
where $y_{ij}\in\{0,0.5,1\}$ corresponds to the conditions that $y_{i}<y_{j},y_{i}=y_{j},y_{i}>y_{j}$, respectively. Following enhancements in the optimization process (Burges et al., 2006) and the incorporation of hinge loss (Tagami et al., 2013) have led to further performance improvements.
Listwise methods encourage positive samples to have higher rankings within the list of all samples. For example, ListNet (Cao et al., 2007) defines its loss as:
(3) $\mathcal{L}_{\text{ListNet}}=-\frac{1}{N_{+}}\sum\limits_{i=1}^{N_{+}}\log\frac{\exp[z_{i}]}{\sum_{k=1}^{N}\exp[z_{k}]}.$
Some other listwise approaches have also achieved remarkable results, such as ListMLE (Xia et al., 2008), BayesRank (Kuo et al., 2009) and PiRank (Swezey et al., 2021).
@@ 2.4. Tailor Ranking Loss into CTR Prediction
In the context of online advertising, recent studies (Sheng et al., 2023; Yan et al., 2022; Bai et al., 2023) argued that it is not easy to achieve decent overall outcomes solely by relying on a single form of the objective function. Consequently, several works that combine classification loss and ranking loss have been proposed with the following objective paradigm:
@@
(5) $\mathcal{L}_{\text{Rank}}^{\text{RCR}}=-\frac{1}{N_{+}}\sum\limits_{i=1}^{N_{+}}\log\frac{\sigma(z_{i})}{\sum_{k=1}^{N}\sigma(z_{k})},$
@@
(6) $\mathcal{L}_{\text{Clf}}^{\text{JRC}}=-\frac{1}{N}\sum\limits_{i=1}^{N}\log\frac{\exp[z_{iy_{i}}]}{\exp[z_{i0}]+\exp[z_{i1}]},$
(7) $\mathcal{L}_{\text{Rank}}^{\text{JRC}}=-\frac{1}{N}\sum\limits_{i=1}^{N}\log\frac{\exp[z_{iy_{i}}]}{\sum_{k=1}^{N}\exp[z_{ky_{i}}]},$
where $z_{i1}$ and $z_{i0}$ are the click/non-click logits of the $i$-th sample, respectively. $z_{ky_{i}}$ is the click logit of $k$-th sample if $i$-th samples is positive. Otherwise, $z_{ky_{i}}$ signifies the non-click logit. In Sec. 3 and Sec. 4, we will conduct detailed empirical experiments and analyses on these methods.
3. Auxiliary Ranking Loss in CTR Prediction: A Gradient Vanishing Perspective
The classification-ranking combination methods are widely used in real-world recommendation systems, e.g., Twitter (Li et al., 2015), Google (Yan et al., 2022; Bai et al., 2023), and Alibaba (Sheng et al., 2023). However, the effectiveness of such methods remains elusive, which drives us to explore the underlying mechanisms at play. In the following, we study Combined-Pair loss (Li et al., 2015) as a representative classification-ranking combination loss and choose DCN V2 (Wang et al., 2021) as the backbone model. The Combined-Pair loss is defined as a combination of a binary cross entropy (BCE) loss and a RankNet loss (Burges et al., 2005):
(8) $\mathcal{L}^{\text{CP}}=\alpha\mathcal{L}_{\text{BCE}}+(1-\alpha)\mathcal{L}_{\text{RankNet}},$
(9) $\mathcal{L}_{\text{RankNet}}=-\frac{1}{N_{+}N_{-}}\sum\limits_{i=1}^{N_{+}}\sum\limits_{j=1}^{N_{-}}\log(\sigma(z_{i}^{(+)}-z_{j}^{(-)})),$
where $N_{+}$ and $N_{-}$ represent the numbers of positive and negative samples, respectively. RankNet within the Combined-Pair is a form of Eq. 2 without $y_{i}=y_{j}$. In the following discussion, we name the model that optimizes the Combined-Pair loss as the Combined-Pair method, or Combined-Pair in short, and the model that optimizes only the BCE loss as the BCE method.
@@ 3.1. Investigation of Classification Ability
Existing research (Li et al., 2015; Sheng et al., 2023) posits that the BCE Loss mainly provides a decent estimate of click probability and involving a ranking loss in addition to the BCE loss improves its ranking performance. However, we are curious about its impact on the main objective of CTR prediction: the classification capability. We train the BCE model and Combined-Pair on the public Criteo dataset (Labs, 2014). Please note that the Criteo dataset’s CTR is 25.6%, which is relatively high due to downsampling on negative samples. A weight $\beta_{\text{pos}}$ for the positive samples was introduced during our model training and evaluation to simulate sparse positive rates in real-world scenarios.
We conducted monitoring of the BCE loss for both the BCE method and the Combined-Pair on the validation set, as depicted by the dashed lines in Fig. 1(a). Interestingly, we made an unexpected observation: the BCE loss of the Combined-Pair method decreases more rapidly than that of the BCE method initially and maintains a consistently lower value thereafter. This finding can be summarized as follows:
Finding 1. Combined-Pair gets a lower BCE loss than the BCE method on the validation set, indicating that it improves the classification ability rather than only the ranking ability.
To investigate whether the improvement in the classification loss is due to better generalization or easier model training when incorporating the ranking loss, we monitored the BCE loss of both methods on the training set, as illustrated by the solid lines in Fig. 1(a). In addition to the previously mentioned surprising results, we made an even more astonishing observation: the BCE loss of the Combined-Pair method on the training set also decreases more rapidly and remains consistently lower than that of the BCE method. We conclude with the second finding:
Finding 2. Combined-Pair gets a lower BCE loss than the BCE method on the training set, indicating that involving an auxiliary ranking loss helps the optimization of the BCE loss.
Besides, we analyze the disparities between the loss landscapes of Combined-Pair and the BCE method. As shown in Fig. 1(b), we observe that the loss landscape of the BCE method exhibits flatter than Combined-Pair.